# FADPNet: Frequency-Aware Dual-Path Network for Face Super-Resolution

### v1

@@ IV-D3 Model Complexity Analysis
Beyond reconstruction quality, we conduct a comprehensive evaluation of model efficiency in terms of parameter count and inference speed, both of which are crucial for devices with limited computing resources. At each stage, inputs are dynamically decomposed into high-frequency and low-frequency components, which are then processed by our specialized sub-networks: a CNN branch tailored for capturing high-frequency facial details, and a Mamba-based branch better suited for modeling low-frequency, long-range dependencies. This innovative design achieves an optimal balance between computational efficiency and an enhanced representational capacity for high-frequency details. By leveraging this structured decomposition, our model adeptly handles complex facial features, leading to improved performance in FSR. As shown in Fig. 11(a) and Fig. 11(b), our method demonstrates an excellent trade-off between reconstruction accuracy and computational cost. Specifically, Fig. 11(a) illustrates that our approach surpass existing methods over 0.2dB PSNR with lower FLOPs and faster inference speed on the CelebA test set, highlighting its practical efficiency. Similarly, Fig. 11(b) further shows that on the Helen test set, our method maintains superior PSNR while keeping the parameter count relatively low, outperforming most existing methods in both accuracy and model compactness. Overall, our approach achieves a well-balanced trade-off between model size, inference speed, and reconstruction quality, making it effective.

### v2

@@ IV-D3 Model Complexity Analysis
Beyond FSR quality, we evaluate model efficiency in terms of parameters and inference speed, critical for resource-constrained devices. At each stage, inputs are decomposed into high- and low-frequency components, processed respectively by a CNN branch for high-frequency facial details and a Mamba-based branch for low-frequency, long-range dependencies. This design balances computational efficiency with strong representation of fine details. As shown in Fig. 10(a) and Fig. 10(b), FADPNet achieves superior PSNR with lower FLOPs, faster inference, and fewer parameters compared to existing methods, consistently outperforming them in both accuracy and model cost.